OpenStudy (anonymous):

evaluate e=0.2, f=4, g=3 (efg)^3

OpenStudy (unklerhaukus):

you could simplify the terms in the parentheses first, or use \[(abc)^n=a^nb^nc^n\]

OpenStudy (anonymous):

idk how to do that

OpenStudy (unklerhaukus):

(efg)^3 = e^3 f^3 g^3

OpenStudy (anonymous):

confused

OpenStudy (unklerhaukus):

(efg)^3 = (efg)(efg)(efg) = eeefffggg = e^3 f^3 g^3

OpenStudy (anonymous):

clueless